Dear Financial Adviser, My spouse and I are each 62 and hope to retire in three years. After retirement we will receive $8,800 per month after taxes from our employers pension plans and $2,800 per month after taxes from Social Security. Unfortunately our monthly living expenses are $16,300. Our social obligations preclude further economies. We have $750,000 invested in a high-grade ta ree munica a bond mutual fund. The return on he fund 1s 4,5% per ear. we an to make annual dra as om hem tua und 0 over hear ere ce between penslon and Soclal Securlty Income and our living expenses. How many years before we run out of money? Sincerely Marblehead, MA You can assume that the withdrawals (one per year) will sit In a checking account (no interest) until spent. The couple will use the account to cover the monthly shortfalls. How many years before Luxury Challenged runs out of money? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Number of years
Answer #1

Answer:

Years to retirement = 3 years

Investment in tax free bond = $750,000

Return of fund = 4.5%

In 3 years fund value = FV = PV * (1 + Rate of return) Number of years

= $750,000 * (1 + 4.5%) 3

= $855,874.59

On retirement:

Monthly expense = $16,300

Monthly after tax inflow from pension and from social security = $8,800 + $2,800 = $11,600

Balance amount required per month = $16,300 - $11,600 = $4,700

Hence, annual withdrawal required = $4,700 * 12 = $56,400 and

Withdrawal has to be at the start of the year

To get number of years (after retirement) before Luxury Challenged will run out of money:

We will use NPER function of excel:

=NPER (rate, pmt, pv, fv, type)

=NPER (4.5%, 56400, -855874.59, 0, 1)

= 24.08 Years

Hence:

Number of years (after retirement) before Luxury Challenged will run out of money = 24.08 years

24.08 Years